First Reply: Future Value & Present Value е Consider the section on Interest (especially in Section 17.3, the sub-section
Answer #1

1.

Given data:

Present value (Left as inheritance) = $100000

Assumed Data:

Interest rate = i = 10%

Years to retire = n = 25 years

Future value = F =?

Since it is not mentioned as to whether we have to take simple interest or compound interest into consideration we shall solve by using both the methods

Simple interest

F = P(1+in)

F = 100000(1+0.10*25)

F = $350000

Compound Interest

F = P(1+i) ^n

F = 100000(1+0.1) ^25

F = $1083470.59

2.

Given data:

Future value = $1000000

Assumed Data:

Interest rate = i = 10%

Years to retire = n = 25 years

Present value = P =?

Since it is not mentioned as to whether we have to take simple interest or compound interest into consideration we shall solve by using both the methods

Simple interest

F = P(1+in)

P = F/(1+in)

P = 1000000 / (1+0.1*25)

P = $285714.3

Compound Interest

F = P(1+i) ^n

P = F / (1+i) ^n

P = 1000000 / (1+0.1) ^25

P = $92296

3. The factors that are held constant are

1. The Interest rate is kept constant, In real life the interest rate tends to fluctuate, however here we consider it to be constant.

2. The inflation is also kept constant or is considered to be zero. If inflation is considered, then the same has to be subtracted from interest rate to find out the effective values.